when four dice are rolled simultaneously, in how many outcomes will at least one of the dice show 3? explain in detail
Answer by scott8148(6628)

possible outcomes __ 4 dice, each with 6 faces __ 6*6*6*6=1296

possible outcomes with no 3 __ 4 dice, only 5 faces allowed __ 5*5*5*5=625

possible outcomes with at least one 3 __ 1296-625=671